OpenStudy (anonymous):

If 12.6 grams of iron (III) oxide reacts with 9.65 grams of carbon monoxide to produce 7.23 g of pure iron, what are the theoretical yield and percent yield of this reaction?

OpenStudy (nikato):

Welcome to openstudy! U actually posted this chemistry question in the math section. http://openstudy.com/study#/groups/Chemistry But I will still help u. 1. Did u find the balanced equation yet?

OpenStudy (anonymous):

Yes, Fe2O3 + 2CO -->2Fe + CO2

OpenStudy (nikato):

Next u want to find the limiting reagent

OpenStudy (nikato):

If u don't know how, check this out http://www.chem.tamu.edu/class/majors/tutorialnotefiles/limiting.htm

OpenStudy (anonymous):

moles Fe2O3 = 12.6 g /159.694 g/mol=0.0789 the ratio between Fe2O3 and CO is 1 : 3 moles CO required = 0.0789 x 3 =0.237 actual moles CO = 9.65 g / 28.01 g/mol=0.345 so CO is in excess the ratio between Fe2O3 and Fe is 1 : 2 theoretical moles Fe = 0.0789 x 2 =0.157 theoretical mass Fe = 0.157 mol x 55.847 g/mol=8.81 % yield = 7.23 x 100/ 8.81 =82.0

OpenStudy (anonymous):

how do you know to divide 12.6 by the molar mass?

OpenStudy (anonymous):

muchas gracias

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS so using this answer will be an academic integrity violation. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065